OpenStudy (poto):

Is this correct? @3mar

OpenStudy (maddielb):

It says it was removed

OpenStudy (steve816):

rip screenshot

OpenStudy (poto):

It should work now http://prntscr.com/d2scgo

OpenStudy (3mar):

Well, I am here.

OpenStudy (poto):

Can you check and see if I got it correct

OpenStudy (steve816):

If a line is perpendicular, than the slope should be the negative reciprocal.

OpenStudy (poto):

So LM and JK

OpenStudy (steve816):

So it would be the opposite of what you chose.

OpenStudy (steve816):

Exactly!

OpenStudy (triciaal):

slope is 4/3 perpendicular slope = -1/m = -1/(4/3) = -3/4

OpenStudy (steve816):

Wait no, it would be LM and GF

OpenStudy (poto):

Okay! thanks everyone!

OpenStudy (triciaal):

choose the segments with slope -3/4

OpenStudy (steve816):

No problem!